OpenStudy (anonymous):

How do I tell if a data set is represented by an exponential function if I have no given f(0)? My reasoning being that if I did, I could use y = Pe^(rt)

OpenStudy (lgbasallote):

you mean for example you are given something like x y 1 1 2 4 3 9 4 16 like that?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

well you could still use y = Pe^(rt) in a way

OpenStudy (lgbasallote):

is this related to exponential growth?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

let's say for example you are given that at t = 2 seconds, a DNA consists of 40 strands. Then, at 4 seconds, the DNA consists of 100 strands. How many DNA strands will there be at 10 seconds? so you have t = 2 and t = 4. You set up the equation like this y = Pe^(rt) \[100 = 40e^{r(4-2)}\] \[100 = 40 e^{2r}\] so now..you can solve for r

OpenStudy (lgbasallote):

does that help?

OpenStudy (lgbasallote):

@jcd2012 still there?

OpenStudy (anonymous):

ya let me look at what you said

OpenStudy (lgbasallote):

tell me when you get it

OpenStudy (anonymous):

so I did what you did for my problem and got r. what next?

OpenStudy (lgbasallote):

in my problem... i asked how many DNA strands there will be in 10 seconds. so i have t = 2 seconds and t = 10 seconds \[y = Pe^{rt}\] \[y = 40e^{r(10 - 2)}\] now you just substitute the value of r there and you can solve for y

OpenStudy (anonymous):

I did that for a known y-value on the data set and the result was not the same

OpenStudy (lgbasallote):

how can you know the y-value of the data if it's not given?

OpenStudy (anonymous):

that's what I'm saying. I'm trying to algebraically show that the data set is from an exponential function when I don't know (0,y)

OpenStudy (lgbasallote):

hmm you want to solve for the value of y at t = 0...let's use the same problem at t = 2, DNA strand is 40 so let's solve how many strands at t = 0 \[y = Pe^{rt}\] \[40 = Pe^{r(2 - 0)}\] substitute r there then you can solve for P

OpenStudy (lgbasallote):

by the way... if you notice, the formula i'm using here is: \[\huge y = Pe^{r(t_2 - t_1)}\] where: y is the final value P is the initial value r is the constant t2 is the final time t1 is the initial time

OpenStudy (anonymous):

still don't have it...

OpenStudy (lgbasallote):

tell me what the value of r is...

OpenStudy (anonymous):

What am I going to get from that? Does it work in your equation? It doesn't in mine

OpenStudy (lgbasallote):

hmm.. let me solve my problem to show you how to do it...

OpenStudy (lgbasallote):

let's say for example you are given that at t = 2 seconds, a DNA consists of 40 strands. Then, at 4 seconds, the DNA consists of 100 strands. How many DNA strands were there initially? \[y = Pe^{rt}\] \[\implies 100 = 40e^{r(4-2)}\] \[\implies 100 = 40 e^{2r}\] \[\implies \frac{100}{40} = e^{2r}\] \[\implies 2.5 = e^{2r}\] \[\implies \ln(2.5) = 2r\] \[\implies 0.9163 = 2r\] \[\implies 0.4582 = r\] now... to solve for P at t = 0 \[y = Pe^{rt}\] \[\implies 40 = Pe^{0.4582(2 - 0)}\] \[\implies 40 = Pe^{0.9163}\] \[\implies 40 = P(2.5)\] \[\implies 16 = P\] so there were 16 strands at t = 0

OpenStudy (lgbasallote):

do you get it now?

OpenStudy (anonymous):

yeah. thanks for sticking it out with me

OpenStudy (lgbasallote):

welcome